OpenStudy (pandaecat):

Check my answer please! Alexei wants to hang a mirror in his boat, and put a frame around it. The mirror and frame must have an area of 19.25 square feet. The mirror is 3 feet wide and 5 feet long. Which quadratic equation can be used to determine the thickness of the frame, x? My answer is 4x^2+16x+15.

OpenStudy (pandaecat):

The choices are 2x^2 + 8x − 19.25 = 0 3x^2 + 5x − 19.25 = 0 4x^2 + 16x − 4.25 = 0 4x^2 + 16x + 15 = 0

OpenStudy (triciaal):